# Tensor Product Attention Is All You Need

### v2

@@ 4.1 Language Modeling Tasks
Training Setup. We follow the nanoGPT training configuration. In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
Training & Validation Curves. Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses. For instance, in Figure 3(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
@@ A.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
$\mathbf{Q}^{(i)}_{t}~{}=~{}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
@@
$\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
@@
$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
@@ A.4 Complexity Analysis for the Specialized Implementation
$\mathbf{Q}^{(i)}_{t}~{}=~{}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
@@
$\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
@@
$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
@@ A.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
Using the factorization $\mathbf{V}_{\tau}=\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
$\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
@@
$\operatorname{head}(t)~{}=~{}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
@@
$\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
@@ A.8 Decoding Speed during Inference Time of MHA, MQA, GQA, MLA, and TPA
$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
@@ G.2.1 Perplexity Curves
We display the perplexity curves for medium, large and XL size of models in Figure 4.
@@ G.2.2 Ablation Study on Different Ranks
Figure 5 shows the training loss, validation loss, and validation perplexity curves of XL-size (1.5B) T6 models with different ranks trained on the FineWeb-Edu 100B dataset, and the evaluation results are displayed in Table 7. It can be observed that increase in rank can improve the performances of large language models.

### v3

@@ 4.1 Language Modeling Tasks
Training Setup. We follow the nanoGPT training configuration. In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $6\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
Training & Validation Curves. Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses. For instance, in Figure 3(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
@@ A.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
@@
$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
@@
$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
@@ A.4 Complexity Analysis for the Specialized Implementation
$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
@@
$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
@@
$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
@@ A.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
Using the factorization $\mathbf{V}_{\tau}=\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
@@
$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
@@
$\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
@@ A.8 Decoding Speed during Inference Time of MHA, MQA, GQA, MLA, and TPA
$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},$
$\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
@@ G.2.1 Perplexity Curves
We display the perplexity curves for medium, large, and XL size models in Figure 4.
@@ G.2.2 Ablation Study on Different Ranks
Figure 5 shows the training loss, validation loss, and validation perplexity curves of XL-size (1.5B) T6 models with different ranks trained on the FineWeb-Edu 100B dataset, and the evaluation results are displayed in Table 7. It can be observed that an increase in rank can improve the performance of large language models.